Prove that the series 1 + (1/2)·(1/3) + [(1·3)/(2· 4)]·(1/5) + [(1·3·5)/(2·4·6)]·(1/7) + ..... converges


1
Expert's answer
2021-02-24T06:39:53-0500

We drop the first term i.e. 1. we consider the rest of the sequence Let us denote each term by "u_n." Then "u_1= \frac{1}{2}\cdot\frac{1}{3}" . We note, "u_n= \frac{1}{2}\cdot\frac{3}{4}\cdots \frac{2n-1}{2n}\frac{1}{2n+1}." Hence "\frac{u_{n}}{u_{n+1}}=\frac{(2n+2)(2n+3)}{(2n+1)^2}."

Hence, "n(\frac{u_{n}}{u_{n+1}}-1)= \frac{6n^2+5n}{4n^2+4n+1}=\frac{6+\frac{5}{n}}{4+\frac{4}{n}+\frac{1}{n^2}}." Hence "lim_{n\rightarrow \infty} \ n(\frac{u_n}{u_{n+1}}-1)=\frac{6}{4}=3/2>1."

Hence by Rabbe's test the series converges.